what is the cube root of x+2 =6

what is the 5th power of x^2 -13 =1

what is the sq.root of 2x + 5 - the sq. root of x-2 = 3


Answer by stanbon(75887)

what is the cube root of x+2 =6
cube both sides to get x+2 = 216
x = 214
-------------------
what is the 5th power of x^2 -13 =1
Take the 5th root of both sides to get:
x^2-13 = 1
x^2 = 14
x = sqrt(14) or x=-sqrt(14)
--------------------------
what is the sq.root of 2x + 5 - the sq. root of x-2 = 3
sqrt(2x+5) - sqrt(x-2) = 3
Square both sides to get:
(2x+5)+(x-2)-2sqrt(2x^2+x-10) = 9
2sqrt(2x^2+x-10) = 3x-6
Square both sides to get:
4(2x^2+x-10) = 9x^2-36x+36
x^2-40x+76 = 0
(x-2)(x-38)=0
x = 2 or x=38
Both answers check
=============================
Cheers,
Stan H.